OpenStudy (anonymous):

What is the equation of the line passing through (–6, 1) and perpendicular to the line y = –3x + 1 in slope-intercept form?

OpenStudy (ivettef365):

first what is the slope on the equation y = -3x + 1

OpenStudy (anonymous):

-3 @ivettef365

OpenStudy (ivettef365):

good, now to be perpendicular it needs to be opposite reciprocal, so what would be the slope

OpenStudy (anonymous):

1/3 i think

OpenStudy (ivettef365):

very good, now the only thing you need to find is the y-int to write the equation of the line

OpenStudy (ivettef365):

so you have y = mx + b ( that is the equation of the line) you have slope = 1/3 points; (-6,1), 1 = 1/3(-6) + b (can you find value of b)

OpenStudy (anonymous):

b=3. so would the equation be y = 1/3x + 3?

OpenStudy (ivettef365):

good, so now you have m = 1/3 and b = 3 so the equation of the line is: y = 1/3m+3

OpenStudy (ivettef365):

sorry I meant y = 1/3x + 3

OpenStudy (anonymous):

okay thank you so much!! i might have more questions later on btw haha

OpenStudy (ivettef365):

yw :)

OpenStudy (anonymous):

What is the equation of a horizontal line passing through the point (–7, 5)